OpenStudy (anonymous):

How to simplify 9x^2+30x+25 / 3x+15 ? Especially how to simplify 9x^2+30x+25???

OpenStudy (unklerhaukus):

you could use the quadratic formula for that numerator, to factorise it, and hopefully one of the factors will cancel

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[(x-x_1)(x-x_2)=0\]

OpenStudy (jhannybean):

Factor out a 3 first from the numerator and denominator.

OpenStudy (anonymous):

thnx Use quadratic make sense.. I dont think we can factor all by 3 (25 doesnt factor 3).....

OpenStudy (jhannybean):

well you know.... 25/3....:P Jk. you're right.

OpenStudy (anonymous):

hehe thats ok! :P

OpenStudy (anonymous):

using quadratic equation: it looks like -30+sqrt(o) / 18 ????

OpenStudy (unklerhaukus):

yes, good stuff, can you simplify further

OpenStudy (anonymous):

yes, -5/3???

OpenStudy (unklerhaukus):

very good., so now you have \[x_{1,2}=-5/3\] what does that look like in this form \[(x-x_1)(x-x_2)=0\]

OpenStudy (anonymous):

I am stuck from here..

OpenStudy (jhannybean):

http://www.sketchtoy.com/42750753

OpenStudy (jhannybean):

An alternative.

OpenStudy (unklerhaukus):

well we are almost there

OpenStudy (anonymous):

is it (x+5/3) and (x-5/3) ??

OpenStudy (unklerhaukus):

there should be no minus sign

OpenStudy (unklerhaukus):

(the factors are the same)

OpenStudy (jhannybean):

\[\large a^2 +b^2 = (a+b)(a+b) = a^2 +2ab+b^2\]

OpenStudy (dan815):

\[\huge a^2 +b^2 = (a+b)(a+b) = a^2 +2ab+b^2\]

OpenStudy (unklerhaukus):

what?

OpenStudy (jhannybean):

\(\large 9x^2 +30x+25\) is a perfect square... \[\large a^2= 9x^2 \ \therefore \ a = 3x\]\[\large b^2 = 25 \ \therefore \ b = 5\]putting this into our formula, \(\large (a+b)(a+b)\) we have \((3x+5)(3x+5)\) or \((3x+5)^2\)

OpenStudy (anonymous):

gotcha! make sense !

OpenStudy (jhannybean):

I was just elaborating what @UnkleRhaukus mentioned earlier....of breaking apart the numerator. He'll continue :P I just thought you could see some other methods.

OpenStudy (dan815):

ooo you made a therefore symbol :O

OpenStudy (dan815):

you are learning new latex symbols

OpenStudy (anonymous):

THnx but I still dont get how to find from here.. (what UnkleRhaukus ) suggest so now you have x1,2=−5/3 what does that look like in this form (x−x1)(x−x2)=0

OpenStudy (unklerhaukus):

x1,2=−5/3 (x+5/3)(x+5/3)=0 (x+5/3)^2 =0

OpenStudy (unklerhaukus):

now multiply both sides of the equation by nine

OpenStudy (anonymous):

(9x+5)^2 for the numerator

OpenStudy (unklerhaukus):

not quite

OpenStudy (unklerhaukus):

the nine becomes a three under the ^2

OpenStudy (anonymous):

9(x+5/3) = (9x+5) so since it is (x+5/3)^2= (9x+5)(9x+5)???

OpenStudy (anonymous):

oh so (3x+5)^2

OpenStudy (unklerhaukus):

\[(x+5/3)^2 =0\\ 9\times(x+5/3)^2 =9\times0\\ 3^2\times(x+5/3)^2 =0\\ (3\times (x+ 5/3))^2=0\\ (3\times x+3\times 5/3)^2=0\]

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

so the result of this equation is (3x+5)^2 / (3x+5) = (3x+5) ..Is that it???

OpenStudy (unklerhaukus):

i thought it was it (9x^2+30x+25) / (3x+15) = (3x+5)^2 / (3x+15)

OpenStudy (unklerhaukus):

does the denominator have +15 or +5?

OpenStudy (anonymous):

darn it!!!!!.. it +5 sry..

OpenStudy (unklerhaukus):

oh well that is good really

OpenStudy (anonymous):

k.. then thank you!!

OpenStudy (unklerhaukus):

(3x+5)^2 / (3x+5) = (3x+5) \(\huge\color{red}\checkmark \)

OpenStudy (unklerhaukus):

@Jhannybean , @dan815 \[\large \color{red}{(a +b)^2} = (a+b)(a+b) = a^2 +2ab+b^2\\ \large\color{red}{a^2 -b^2=(a+b)(a-b)=a^2-ab+ba-b^2}\\ \large\color{brown}{a^2 +b^2=(a-ib)(a+ib)=a^2+aib-iba+b^2}\]

OpenStudy (anonymous):

thats really helps!!

OpenStudy (dan815):

\[\large \color{navy}{(a +b)^2 =(a+b)(a+b) = a^2 +2ab+b^2}\\\] \[\large\color{blue}{a^2 -b^2=(a+b)(a-b)=a^2-ab+ba-b^2}\\\] \[\large\color{skyblue}{a^2 +b^2=(a-ib)(a+ib)=a^2+aib-iba+b^2}\]

OpenStudy (dan815):

they shud have some crazy crayola color names put in too

OpenStudy (dan815):

red mountain glaze

OpenStudy (unklerhaukus):

you can use hex colours http://openstudy.com/study#/updates/51b0e7f9e4b05b167ed2e8fe

OpenStudy (jhannybean):

Ahh... Then I was wrong.